### v1

# Efficient Diffusion Model for Image Restoration by Residual Shifting

Abstract
While diffusion-based image restoration (IR) methods have achieved remarkable success, they are still limited by the low inference speed attributed to the necessity of executing hundreds or even thousands of sampling steps. Existing acceleration sampling techniques, though seeking to expedite the process, inevitably sacrifice performance to some extent, resulting in over-blurry restored outcomes. To address this issue, this study proposes a novel and efficient diffusion model for IR that significantly reduces the required number of diffusion steps. Our method avoids the need for post-acceleration during inference, thereby avoiding the associated performance deterioration. Specifically, our proposed method establishes a Markov chain that facilitates the transitions between the high-quality and low-quality images by shifting their residuals, substantially improving the transition efficiency. A carefully formulated noise schedule is devised to flexibly control the shifting speed and the noise strength during the diffusion process. Extensive experimental evaluations demonstrate that the proposed method achieves superior or comparable performance to current state-of-the-art methods on three classical IR tasks, namely image super-resolution, image inpainting, and blind face restoration, even only with four sampling steps. Our code and model are publicly available at https://github.com/zsyOAOA/ResShift.
Index Terms:
Markov chain, noise schedule, image super-resolution, image inpainting, face restoration.
I Introduction
Image restoration (IR) is a fundamental problem in low-level vision, aiming at recovering a high-quality (HQ) image given its corresponding low-quality (LQ) counterpart. It encompasses various sub-tasks upon the type of degradation, including but not limited to image denoising [1, 2], image super-resolution [3, 4] and image inpainting [5, 6]. Notably, the degradation model is usually complicated in some cases, e.g., real-world super-resolution, making the IR problem severely ill-posed and challenging.
The diffusion model [7], a newly emerged generative model, has demonstrated remarkable success in the realm of image generation [8, 9], surpassing the previous GAN-based approaches [10, 11]. The main idea of diffusion models is to construct a hidden Markov chain that gradually degrades an image to white Gaussian noise in the forward process, and then approximate the reverse process using a deep neural network to generate this image. Attributed to its powerful generative capability, diffusion model has also exhibited great promise in solving the IR tasks, including image denoising [12, 13], deblurring [14, 15], inpainting [16, 17, 18, 19], colorization [20, 21, 22], among others. The exploration of the potential of diffusion models in IR still remains an active and ongoing research pursuit.
Existing diffusion-based IR methods can be broadly categorized into two classes. One straightforward approach [23, 24, 21, 25, 26, 27] is to insert the LQ image into the input of current diffusion model, e.g., DDPM [8], as a condition. Then, the model is retrained specifically for the task of IR. After training, this model is able to generate a desirable HQ image from Gaussian noise and the observed LQ image via the reverse sampling process. Another popular way, as explored in [28, 29, 30, 31, 32, 33, 34], attempts to harness a pre-trained unconditional diffusion model as a prior to facilitate the IR problem. The primary idea involves modifying the reverse sampling procedure to align the generated result with the given LQ observation via the degradation model at each step. Unfortunately, both strategies inherit the Markov chain underlying DDPM, which can be inefficient during inference, i.e., often necessitating hundreds or even thousands of sampling steps. Although some acceleration techniques [35, 36, 37] have been developed to reduce the sampling steps in inference, they inevitably lead to a significant drop in performance, resulting in over-smooth results as shown in Fig. 1(e)-(j), in which the DDIM [36] algorithm is employed to speed up the inference. Therefore, there is a need to devise a new diffusion model tailored for IR, capable of achieving a harmonious balance between efficiency and performance, without sacrificing one for the other.
Let us revisit the diffusion model within the context of image generation. In the forward process, it constructs a Markov chain to gradually transform the observed data into a pre-specified prior distribution, typically a standard Gaussian distribution, over a large number of steps. Conversely, the reverse process trains a deep neural network to approximate the reverse trajectory of this Markov chain. The well-trained neural network facilitates the random image generation by drawing samples from the reverse Markov chain, initiating at the Gaussian distribution. While the Gaussian distribution is well-suited for the task of image generation, its optimality is questioned in the domain of IR, where the LQ images are available. In this paper, we argue that a reasonable diffusion model for IR should start from a prior distribution centered around the LQ image, enabling an iterative recovery of the HQ image from its LQ counterpart instead of Gaussian white noise. Notably, this design not only aligns with the specifics of IR but also holds the potential to reduce the requisite number of diffusion steps for sampling, thereby enhancing overall inference efficiency.
Figure 1: Qualitative comparisons on one typical real-world example of the proposed method and recent state of the arts, including RealESRGAN [38], BSRGAN [39], SwinIR [40], LDM [25], and StableSR [29]. As for the diffusion-based approaches, namely LDM, StableSR, and our proposed method, we annotate the number of sampling steps with the format of “Method-A” for more intuitive visualization, where “A” denotes the number of sampling steps. Note that LDM and StableSR both contain 1000 diffusion steps in training and are accelerated to “A” steps using DDIM [36] during inference. Please zoom in for a better view.
Following the aforementioned motivation, we propose an efficient diffusion model involving a shorter Markov chain for the transition between the HQ image and its corresponding LQ one. The initial state of the Markov chain converges towards an approximate distribution of the HQ image, while the final state converges towards an approximate distribution of the LQ image. To achieve this, we carefully design a transition kernel that gradually shifts the residual information between the HQ/LQ image pair. This approach exhibits superior efficiency in comparison to existing diffusion-based IR methods, given its ability to expeditiously transfer the residual information in several steps. Moreover, our design also allows for an analytical and concise expression for the evidence lower bound, thereby simplifying the formulation of the optimization objective for training purposes. Building upon the constructed diffusion kernel, we further develop a highly flexible noise schedule that controls the speed of residual shifting and the strength of the added noise in each step. This schedule facilitates a fidelity-realism trade-off of the recovered results by tuning its hyper-parameters.
In summary, the main contributions of this work are as follows:
  • We propose an efficient diffusion model specifically for IR. It builds up a short Markov chain between the HQ/LQ images, rendering a fast reverse sampling trajectory during inference. Extensive experiments show that our approach requires only four sampling steps to achieve appealing results, outperforming or at least being comparable to current state-of-the-art methods. A preview of the comparison results of the proposed method to recent approaches is shown in Fig. 1.
  • A highly flexible noise schedule is designed for the proposed diffusion model, capable of controlling the transition properties more precisely, including the shifting speed and the noise level. Through tuning the hyper-parameters, our method offers a more graceful solution to the widely acknowledged perception-distortion trade-off in IR.
  • The proposed method is a general diffusion-based framework for IR, being able to handle various IR tasks. This study has thoroughly substantiated its effectiveness and superiority on three typical and challenging IR tasks, namely image super-resolution, image inpainting, and blind face restoration.
In summary, this work endeavors to formulate an efficient diffusion model customized for IR, with the intention of breaking down the limitation of prevailing approaches on inference efficiency. A preliminary version of this work has been published in NeurIPS 2023 [41], focusing only on the task of image super-resolution. This study makes substantial improvements in both model design and empirical evaluation across diverse IR tasks compared with the conference version. Concretely, we incorporate the perceptual loss into the model optimization and substitute the self-attention layer with Swin transformer in the network architecture. The former modification can further reduce the diffusion steps from 15 to 4, and the latter endows our model with graceful adaptability to handle arbitrary resolutions during inference.
The remainder of the manuscript is organized as follows: Sec. II introduces the related work. Sec. III presents our designed diffusion model for IR. In Sec. IV and Sec. V, extensive experiments are conducted to evaluate the performance of our proposed method on the task of image super-resolution and image inpainting, respectively. Sec. VI finally concludes the paper.
II Related Work
In this section, we briefly review the literature on image restoration, traversing from conventional non-diffusion methodologies to recent diffusion-based approaches.
II-A Conventional Image Restoration Approaches
Most of the conventional IR methods can be cast into the Maximum a Posteriori (MAP) framework, a Bayesian paradigm encompassing a likelihood (loss) term and a prior (regularization) term. The likelihood reflects the underlying noise distribution of the LQ image. The commonly used $L_{2}$ or $L_{1}$ loss indeed corresponds to a Gaussian or Laplacian assumption on image noise. To more accurately depict the noise distribution, some robust alternatives were introduced, such as Poissonian-Gaussian [42], MoG [43], MoEP [44], Dirichlet MoG [45, 46] and so on. Simultaneously, there has been an increased focus on employing image priors to address the inherent ill-posedness of IR over recent decades. Typical image priors encompass total variation [47], wavelet coring [48], non-local similarity [49, 3], sparsity [50, 51], low-rankness [1, 52], dark channel [53, 54], among others. These conventional methods are mainly limited by the model capacity and the subjectivity inherited from the manually designed assumptions on image noise and prior.
In recent years, the landscape of IR has been dominated by deep learning (DL)-based methodologies. The seminal works [4, 55, 2] proposed to solve the IR problem using a convolution neural network, outperforming traditional model-based methods on the tasks of image denoising, super-resolution, and deblurring. Then, many studies [56, 57, 58, 59, 60, 61, 62, 63, 64, 6] have emerged, mainly concentrating on designing more delicate network architectures to further improve the restoration performance. Besides, there have been some discernible investigations that seek to combine current DL tools and classical IR modeling ideas. Notable works include the plug-and play or unfolding paradigm [65, 66, 67], learnable image priors [68, 69, 70, 71], and the loss-oriented methods [72, 73, 74]. The infusion of deep neural networks, owing to their large model capacity, has substantively extended the frontiers of IR tasks.
II-B Diffusion-based Image Restoration Approaches
Inspired by principles from non-equilibrium statistical physics, Sohl-Dickstein et al. [7] proposed the diffusion model to fit complex distributions. Subsequent advancements by Ho et al. [8] and Song et al. [20] further improve its theoretical foundation by integrating denoising score matching and stochastic differential equation, thereby achieving impressive success in image generation [25]. Owing to its powerful generative capability, diffusion models have also found successful applications in the field of IR. Next, we provide a comprehensive overview of recent developments in diffusion-based IR methods.
The most straightforward solution to solve the IR problem using diffusion models is to introduce the LQ image as an addition condition in each timestep. Pioneering this approach, Saharia et al. [23] have successfully trained a diffusion model for image super-resolution. Subsequent studies [21, 14, 27] further expanded upon this approach, exploring its applicability in image deblurring, colorization, and denoising. To circumvent the resource-intensive process of training from scratch, an alternative strategy involves harnessing a pre-trained diffusion model to facilitate IR tasks. Numerous investigations, such as [18, 24, 28, 17, 32, 33, 15], reformulated the reverse sampling procedure of a pre-trained diffusion model into an optimization problem by incorporating the degradation model, enabling solving the IR problem in a zero-shot manner. Most of these methods, however, cannot deal with the blind IR problem, as they rely on a pre-defined degradation model. In contrast, some other works [29, 75, 30] directly introduced a trainable module that takes the LQ image as input. This module modulates the feature maps of the pre-trained diffusion model, steering it toward the direction of generating a desirable HQ image. Such a paradigm eliminates the reliance on a degradation model in the test phase, rendering it capable of handling the blind IR tasks.
The methodologies mentioned above are grounded in the foundational diffusion model initially crafted for image generation, necessitating a large number of sampling steps. This inefficiency presents a constraint on their application in real scenarios. The primary goal of our investigation is to devise a new diffusion model customized for IR, which facilitates a swift transition between the LQ/HQ image pair, thereby enhancing efficiency during inference.
III Methodology
In this section, we present our proposed diffusion model tailored for IR. For ease of presentation, the LQ image and the corresponding HQ image are denoted as $\bm{y}_{0}$ and $\bm{x}_{0}$, respectively. Notably, we further assume $\bm{y}_{0}$ and $\bm{x}_{0}$ have identical spatial resolution, which can be easily achieved by pre-upsampling the LQ image $\bm{y}_{0}$ using nearest neighbor interpolation if necessary.
Figure 2: Overview of the proposed method. Our method builds up a Markov chain between the HQ/LQ image pair by shifting their residuals. To alleviate the computational burden of this transition, it can be optionally moved to the latent space of VQGAN [76].
III-A Model Design
The iterative sampling paradigm of diffusion models has proven highly effective in generating intricate and vivid image details, inspiring us to design an iterative approach to address the IR problem as well. Our proposed method builds up a Markov chain to facilitate a transition from the HQ image to its LQ counterpart as shown in Fig. 2. In this way, the restoration of the desirable HQ image can be achieved by sampling along the reverse trajectory of this Markov chain that starts at the given LQ image. Next, we will detail how to construct such a Markov chain specifically for IR.
Forward Process. Let’s denote the residual between the LQ image and its corresponding HQ counterpart as $\bm{e}_{0}$, i.e., $\bm{e}_{0}=\bm{y}_{0}-\bm{x}_{0}$. Our core idea is to construct a transition from $\bm{x}_{0}$ to $\bm{y}_{0}$ by gradually shifting their residual $\bm{e}_{0}$ through a Markov chain with length $T$. Before that, we first introduce a shifting sequence $\{\eta_{t}\}_{t=1}^{T}$, which increases monotonically with respect to timestep $t$ and adheres to the condition of $\eta_{1}\to 0$ and $\eta_{T}\to 1$. Then, the transition distribution is formulated based on this shifting sequence as follows:
$q(\bm{x}_{t}|\bm{x}_{t-1},\bm{y}_{0})=\mathcal{N}(\bm{x}_{t};\bm{x}_{t-1}+\alpha_{t}\bm{e}_{0},\kappa^{2}\alpha_{t}\bm{I}),~{}t=1,\cdots,T,$ (1)
where $\alpha_{1}=\eta_{1}$ and $\alpha_{t}=\eta_{t}-\eta_{t-1}$ for $t>1$, $\kappa$ is a hyper-parameter controlling the noise variance, $\bm{I}$ is the identity matrix. Notably, we show that the marginal distribution at any timestep $t$ is analytically integrable, namely
$q(\bm{x}_{t}|\bm{x}_{0},\bm{y}_{0})=\mathcal{N}(\bm{x}_{t};\bm{x}_{0}+\eta_{t}\bm{e}_{0},\kappa^{2}\eta_{t}\bm{I}),~{}t=1,\cdots,T.$ (2)
The design of the transition distribution presented in Eq. (1) is guided by two primary principles. The first principle concerns the standard deviation, i.e., $\kappa\sqrt{\alpha_{t}}$, which aims to facilitate a smooth transition between $\bm{x}_{t}$ and $\bm{x}_{t-1}$. This is achieved by bounding the expected distance between $\bm{x}_{t}$ and $\bm{x}_{t-1}$ with $\sqrt{\alpha_{t}}$, given that the image data falls within the range of $[0,1]$. Mathematically, this is expressed as:
$\text{max}[(\bm{x}_{0}+\eta_{t}\bm{e}_{0})-(\bm{x}_{0}+\eta_{t-1}\bm{e}_{0})]=\text{max}[\alpha_{t}\bm{e}_{0}]<\alpha_{t}<\sqrt{\alpha_{t}},$ (3)
[…]
Reverse Process. The reverse process endeavors to estimate the posterior distribution $p(\bm{x}_{0}|\bm{y}_{0})$ through the following formalization:
$p(\bm{x}_{0}|\bm{y}_{0})=\int p(\bm{x}_{T}|\bm{y}_{0})\prod_{t=1}^{T}p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{y}_{0})\mathrm{d}\bm{x}_{1:T},$ (4)
where $p(\bm{x}_{T}|\bm{y}_{0})\approx\mathcal{N}(\bm{x}_{T}|\bm{y}_{0},\kappa^{2}\bm{I})$, $p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{y}_{0})$ represents the desirable inverse transition kernel from $\bm{x}_{t}$ to $\bm{x}_{t-1}$, parameterized by a learnable parameter $\bm{\theta}$. Consistent with prevalent literature of diffusion model [7, 8, 20], we adopt the following Gaussian assumption:
$p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{y}_{0})=\mathcal{N}\left(\bm{x}_{t-1};\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t),\bm{\Sigma}_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)\right).$ (5)
The optimization for $\bm{\theta}$ is achieved by minimizing the negative evidence lower bound, namely,
$\min_{\bm{\theta}}\sum_{t}D_{\text{KL}}\left[q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0},\bm{y}_{0})\|p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{y}_{0})\right],$ (6)
where $D_{\text{KL}}[\cdot\|\cdot]$ denotes the Kullback-Leibler (KL) divergence. More mathematical details can be found in [7] or [8].
By combining Eq. (1) and Eq. (2), the target distribution $q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0},\bm{y}_{0})$ in Eq. (6) can be rendered tractable and expressed in an explicit form given below:
$q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0},\bm{y}_{0})=\mathcal{N}\left(\bm{x}_{t-1}\bigg{|}\frac{\eta_{t-1}}{\eta_{t}}\bm{x}_{t}+\frac{\alpha_{t}}{\eta_{t}}\bm{x}_{0},\kappa^{2}\frac{\eta_{t-1}}{\eta_{t}}\alpha_{t}\bm{I}\right).$ (7)
The detailed calculation of this derivation can be found in Appendix -A. Considering that the variance parameter is independent of $\bm{x}_{t}$ and $\bm{y}_{0}$, we thus set it to be the true variance, i.e.,
$\bm{\Sigma}_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)=\kappa^{2}\frac{\eta_{t-1}}{\eta_{t}}\alpha_{t}\bm{I}.$ (8)
The mean parameter $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)$ is reparameterized as:
$\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)=\frac{\eta_{t-1}}{\eta_{t}}\bm{x}_{t}+\frac{\alpha_{t}}{\eta_{t}}f_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t),$ (9)
where $f_{\bm{\theta}}(\cdot)$ is a deep neural network with parameter $\bm{\theta}$, aiming to predict $\bm{x}_{0}$. We explored different parameterization forms on $\bm{\mu}_{\theta}$ and found that Eq. (9) exhibits superior stability and performance.
Based on Eq. (9), the objective function in Eq. (6) is then simplified as:
$\min_{\bm{\theta}}\sum\nolimits_{t}w_{t}\|f_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)-\bm{x}_{0}\|_{2}^{2},$ (10)
where $w_{t}=\frac{\alpha_{t}}{2\kappa^{2}\eta_{t}\eta_{t-1}}$. In practice, we empirically find that the omission of weight $w_{t}$ results in a notable performance improvement, aligning with the conclusion in [8].
Perceptual Regularization. As presented above, our proposed method facilitates an iterative restoration process starting from the LQ image, in contrast to prior methods that initialize from Gaussian noise. This approach effectively reduces the number of diffusion steps. A comprehensive experimental analysis in Sec. IV-B substantiates that our method yields promising results with a mere 15 sampling steps, demonstrating a notable acceleration compared to established methodologies [25, 29].
Unfortunately, attempts at further acceleration, particularly with fewer than 5 sampling steps, tend to produce over-smooth results. This phenomenon is primarily attributed to that the $L_{2}$-based loss in Eq. (10) favors the prediction of an average over plausible solutions. To overcome this limitation, we introduce the perceptual regularization [77] to impose an additional constraint on the solution space, namely,
$\min_{\bm{\theta}}\sum\nolimits_{t}\|f_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)-\bm{x}_{0}\|_{2}^{2}+\lambda l_{p}\left(f_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t),\bm{x}_{0}\right),$ (11)
where $l_{p}(\cdot,\cdot)$ denotes the pre-trained LPIPS metric, $\lambda$ is a hyper-parameter controlling the relative importance of these two constraints. This simple solution effectively curtails the sampling trajectory to fewer steps, e.g., 4 steps in this study, as shown in Sec. IV-B, while concurrently maintaining superior performance.
Extension to Latent Space. To alleviate the computational overhead in training, our proposed model can be optionally moved into the latent space of VQGAN [76], where the original image is compressed by a factor of four in spatial dimensions. This does not require any modifications on our model other than substituting $\bm{x_{0}}$ and $\bm{y}_{0}$ with their latent codes. An intuitive illustration is shown in Fig. 2.
Figure 3: Illustration of the proposed noise schedule. (a) HQ image. (b) Zoomed LQ image. (c)-(d) Diffused images of the proposed noise schedule in timesteps of 1, 3, 5, 7, 9, 12, and 15 under different values of $\kappa$ by fixing $p=0.3$ and $T=15$. (e)-(f) Diffused images of our method with a specified configuration of $\kappa=40,p=0.8,T=1000$ and LDM [25] in timesteps of 100, 200, 400, 600, 800, 900, and 1000. (g) The relative noise intensity (vertical axes, measured by $\sqrt{\nicefrac{{1}}{{\lambda_{\text{snr}}}}}$, where $\lambda_{\text{snr}}$ denotes the signal-to-noise ratio) of the schedules in (d) and (e) w.r.t. the timesteps (horizontal axes). (h) The shifting speed $\sqrt{\eta_{t}}$ (vertical axes) w.r.t. to the timesteps (horizontal axes) across various configurations of $p$. Note that the diffusion processes in this figure are implemented in the latent space, but we display the intermediate results after decoding back to the image space for the purpose of easy visualization.
III-B Noise Schedule
The proposed method employs a hyper-parameter $\kappa$ and a shifting sequence $\{\eta_{t}\}_{t=1}^{T}$ to determine the noise schedule of the diffusion process. In particular, the hyper-parameter $\kappa$ regulates the overall noise intensity during the transition, and its influence on performance is empirically discussed in Sec. IV-B. The subsequent exposition mainly revolves around the construction of the shifting sequence $\{\eta_{t}\}_{t=1}^{T}$.
Equation (2) indicates that the stochastic perturbation in state $\bm{x}_{t}$ is proportional to $\sqrt{\eta_{t}}$, incorporating a scaling factor $\kappa$. This observation motivates us to focus on the design of $\sqrt{\eta_{t}}$ rather than $\eta_{t}$. Previous work by Song et al. [78] has suggested that maintaining a sufficiently small value for $\kappa\sqrt{\eta_{1}}$, such as 0.04 in LDM [25], is imperative to ensure $q(\bm{x}_{1}|\bm{x}_{0},\bm{y}_{0})\approx q(\bm{x}_{0})$. Further considering $\eta_{1}\to 0$, we set $\eta_{1}$ to be the minimum value between $(\nicefrac{{0.04}}{{\kappa}})^{2}$ and $0.001$. For the terminal step $T$, we set $\eta_{T}$ as 0.999, guaranteeing $\eta_{T}\to 1$. For the intermediate timesteps $t\in[2,T-1]$, we propose a non-uniform geometric schedule for $\sqrt{\eta_{t}}$ as follows:
$\sqrt{\eta_{t}}=\sqrt{\eta_{1}}\times b_{0}^{\beta_{t}},~{}t=2,\cdots,T-1,$ (12)
where
$\beta_{t}$ $=\left(\frac{t-1}{T-1}\right)^{p}\times(T-1),$ (13a)
$b_{0}$ $=\exp\left[\frac{1}{2(T-1)}\log{\frac{\eta_{T}}{\eta_{1}}}\right].$ (13b)
It should be noted that the choice of $\beta_{t}$ and $b_{0}$ is grounded in the assumption of $\beta_{1}=0$, $\beta_{T}=T-1$, and $\sqrt{\eta_{T}}=\sqrt{\eta_{1}}\times b_{0}^{T-1}$. The hyper-parameter $p$ controls the growth rate of $\sqrt{\eta_{t}}$, as depicted in Fig. 3(h).
The proposed noise schedule exhibits a high degree of flexibility in three key aspects. First, in the case of small values of $\kappa$, the final state $\bm{x}_{T}$ converges towards a perturbation around the LR image, as illustrated in Fig. 3(c)-(d). Compared to the diffusion process ended at Gaussian noise, this design significantly shortens the length of the Markov chain, thereby improving the inference efficiency. Second, the hyper-parameter $p$ provides precise control over the shifting speed, enabling a fidelity-realism trade-off in the SR results as analyzed in Sec. IV-B. Third, by setting $\kappa=40$ and $p=0.8$, our method achieves a diffusion process that remarkably resembles to LDM [25]. This is clearly demonstrated by the visual results of the diffusion process presented in Fig. 3(e)-(f), and further supported by the comparisons on the relative noise strength as shown in Fig. 3(g).
Figure 4: Visual comparison of two different models containing some self-attention layers (denoted as model-1) or Swin transformers (denoted as model-2). (a1) and (a2): zoomed LQ images with resolutions of $64\times 64$ or $128\times 128$. (b1) and (b2): super-resolved results by model-1. (c1) and (c2): visualized attention maps extracted from the first self-attention layer of model-1. Note that these visualized results are obtained by first calculating the first principal component of the attention map via PCA and then reshaping it to the targeted size. In the left-upper corner, we annotate the average entropy values of these attention maps. (d1) and (d2): super-resolved results by model-2.
@articleliang2024control, title=Control Color: Multimodal Diffusion-based Interactive Image Colorization, author=Liang, Zhexin and Li, Zhaochen and Zhou, Shangchen and Li, Chongyi and Loy, Chen Change, journal=arXiv preprint arXiv:2402.10855, year=2024
III-C Discussion on Arbitrary Resolution
It is widely acknowledged that the self-attention layer [79], a pivotal component in recent diffusion architectures, plays a crucial role in capturing global information in image generation. In the field of IR, however, it causes a blurring issue in handling the test images with arbitrary resolutions, particularly when the test resolution largely diverges from the training resolution. One typical example is provided in Figure 4, considering two LQ images with different resolutions. The baseline model with multiple self-attention layers, which is trained on a resolution of $64\times 64$, performs well when the LQ image aligns with the training resolution but yields blurred results when confronted with a mismatched resolution of $128\times 128$.
To analyze the underlying reason, we visualize the attention maps extracted from the first attention layer in this baseline network, as shown in Fig. 4(c1) and (c2). Note that these two attention maps are both interpolated to the resolution of $256\times 256$ for ease of comparison. Evidently, the example with a larger resolution tends to generate a more uniformly distributed attention map, i.e., Fig. 4(c2), being consistent with the entropy values annotated on the left-upper corner [^2] The principle of maximum entropy posits that it achieves the maximum entropy when the attention map conforms to a uniform distribution. Consequently, a uniformly distributed attention map often leads to an over-smooth outcome, introducing undesirable distortions in performance.
To address this issue, some recent studies [14, 12] have chosen to discard the self-attention layers, a strategy that typically results in a noticeable decline in performance. Inspired by Liang et al. [40], we propose a solution by substituting the self-attention layer with that in Swin Transformer [80]. This straightforward replacement not only alleviates the blurring problem but also maintains the promised performance, as shown in Fig. 4(d1) and (d2). This is because the Swin Transformer computes the attention map in a local window, thus being independent of the image resolution.
IV Experiments on Image Super-resolution
This section offers an evaluation of the proposed method on the task of image super-resolution (SR), with a particular focus on the setting of ×$\times 4$ SR following existing studies [39, 38]. We first provide an in-depth analysis on the proposed model and then conduct a thorough comparison against recent state-of-the-art methods (SotAs) on one synthetic and two real-world datasets. For the sake of brevity in presentation, our method is herein referred to ResShift or ResShiftL. The former is trained based on the primary loss of diffusion model in Eq. (10), while the latter further introduces the perceptual regularization as shown in Eq. (11). To facilitate a more profound understanding of our designed diffusion model and the noise schedule, the model analysis part of Sec. IV-B mainly centers around ResShift.
Figure 5: Qualitative comparisons of ResShift under different combinations of ($T$, $p$, $\kappa$) on the task of image super-resolution. For example, “(15, 0.3, 2.0)” represents the recovered result with $T=15$, $p=0.3$, and $\kappa=2.0$. Please zoom in for a better view.
TABLE I: Performance comparison of ResShift on the ImageNet-Test dataset for image super-resolution under different configurations.
  Configurations Metrics
$T$
$p$
$\kappa$
[…]
4
0.3 2.0
25.64
0.6903
0.3242
10
25.20
0.6828
0.2517
15
25.01
0.6769
0.2312
30
24.52
0.6585
0.2253
50
24.22
0.6483
0.2212
15
0.3
2.0
25.01
0.6769
0.2312
0.5
25.05
0.6745
0.2387
1.0
25.12
0.6780
0.2613
2.0
25.32
0.6827
0.3050
3.0
25.39
0.5813
0.3432
15 0.3
0.5
24.90
0.6709
0.2437
1.0
24.84
0.6699
0.2354
2.0
25.01
0.6769
0.2312
8.0
25.31
0.6858
0.2592
16.0
24.46
0.6891
0.2772
IV-A Experimental Setup
[…]
Figure 6: Perception-distortion trade-off of ResShift and LDM. The vertical and horizontal axes represent the strength of the perception and distortion, measured by LPIPS and MSE, respectively.
Test Datasets. We randomly select 3000 images from the validation set of ImageNet [81] as our synthetic test data, denoted as ImageNet-Test for convenience. The LQ images are generated based on the commonly-used degradation model:
$\bm{y}=(\bm{x}*\bm{k})\downarrow+\bm{n},$ (14)
where $\bm{k}$ is the blurring kernel, $\bm{n}$ is the noise, $\bm{y}$ and $\bm{x}$ denote the LQ image and HQ image, respectively. To comprehensively evaluate the performance of our model, we consider more complicated types of blurring kernel, downsampling operator, and noise type. The detailed settings on them can be found in Appendix -C1. It should be noted that we select the HQ images from ImageNet [81] instead of the prevailing datasets in SR, such as Set5 [85], Set14 [86], and Urban100 [87]. The rationale behind this setting is rooted in the fact that these prevailing datasets only contain very few source images, which fail to thoroughly evaluate the performance of various methods under a multitude of different degradation types.
Two real-world datasets are adopted to evaluate the efficacy of our method. The first is RealSR-V3 [88], containing 100 real images captured by Canon 5D3 and Nikon D810 cameras. Additionally, we collect another real-world dataset named RealSet80. It comprises 50 LR images widely used in recent literature [89, 90, 91, 92, 38, 93]. The remaining 30 images are downloaded from the internet by ourselves.
[…]
Methods
Metrics
6
[…]
Runtime (s)
#Params (M)
ESRGAN [72]
20.67
0.448
0.485
0.038
16.70
RealSR-JPEG [94]
23.11
0.591
0.326
0.038
16.70
BSRGAN [39]
24.42
0.659
0.259
0.038
16.70
SwinIR [40]
23.99
0.667
0.238
0.107
28.01
RealESRGAN [38]
24.04
0.665
0.254
0.038
16.70
DASR [95]
24.75
0.675
0.250
0.022
8.06
LDM-15 [25]
24.89
0.670
0.269
0.247
[…]
LDM-4 [25]
24.74
0.657
0.345
0.077
StableSR-15 [29]
23.37
0.631
0.262
1.070
[…]
StableSR-4 [29]
24.11
0.658
0.287
0.399
ResShift-15
25.01 0.677 0.231
0.682
[…]
ResShiftL-4
25.02 0.683 0.208
0.186
Figure 7: Qualitative results of different methods on the synthetic ImageNet-Test dataset for image super-resolution. Note that we only display the comparison results to the recent five SotA methods in (b)-(f) due to the page limitation, and the whole results can be found in Fig. 14 of the appendix. Please zoom in for a better view.
TABLE III: Quantitative results of different methods on two real-world datasets for image super-resolution. Note that the results of diffusion-based methods are denoted as “Method-A”, where “A” represents the number of sampling steps. The best and second best results are highlighted in bold and underline.
Methods
Datasets
5
RealSR-V3 [88] RealSet80
5
[…]
ESRGAN [72]
0.2362
29.048
0.4165
48.153
RealSR-JPEG [94]
0.3615
36.076
0.5828
57.379
BSRGAN [39]
0.5439
63.586
0.6263
66.629
SwinIR [40]
0.4654
59.632
0.6072
64.739
RealESRGAN [38]
0.4898
59.678
0.6189
64.496
DASR [95]
0.3629
45.825
0.5311
58.974
LDM-50 [25]
0.4907
54.391
0.5511
55.826
LDM-15 [25]
0.3836
49.317
0.4592
50.972
StableSR-50 [29]
0.5208
60.177
0.6214
62.761
StableSR-15 [29]
0.4974
59.099
0.5975
61.476
ResShift-15
0.6014
58.979
0.6645
62.782
ResShiftL-4 0.5854
56.184
0.6483
61.017
Compared Methods. We evaluate the effectiveness of ResShift and ResShiftL in comparison to seven recent SR methods, namely RealSR-JPEG [94], BSRGAN [39], RealESRGAN [38], SwinIR [40], DASR [95], LDM [25], and StableSR [29]. LDM and StableSR are both diffusion-based methods trained with 1,000 diffusion steps. For a fair comparison, we accelerate them to 15 or 4 steps using the DDIM [36] algorithm during inference in this study. The hyper-parameter $\eta$ in DDIM is set to be 1.0 as this value yields the most realistic results. For clarity, the results of LDM and StableSR are denoted as “Method-A”, where “A” represents the number of inference steps.
Evaluation Metrics. We evaluate the efficacy of different methods using three widely used metrics in most of the SR literature, namely PSNR, SSIM [96], and LPIPS [77], on the synthetic dataset. Notably, PSNR and SSIM are calculated in the luminance (Y) channel of YCbCr space, while LPIPS is directly computed in the standard RGB (sRGB) space. For assessments on real-world datasets, we mainly rely on CLIPIQA [97] and MUSIQ [98] as evaluative metrics to compare the performance of various methods. CLIPIQA and MUSIQ are both non-reference metrics specifically designed for assessing the realism of images. Of particular note, CLIPIQA leverages the CLIP [99] model, pre-trained on the extensive Laion400M [100] dataset, thereby demonstrating strong generalization ability.